# Calabi-Yau Four/Five/Six-folds as $\mathbb{P}_{\textbf{w}}^{n}$ Hypersurfaces: Machine Learning, Approximation, and Generation

### v1

@@ Abstract
Calabi-Yau four-folds may be constructed as hypersurfaces in weighted projective spaces of complex dimension 5 defined via weight systems of 6 weights. In this work, neural networks were implemented to learn the Calabi-Yau Hodge numbers from the weight systems, where gradient saliency and symbolic regression then inspired a truncation of the Landau-Ginzburg model formula for the Hodge numbers of any dimensional Calabi-Yau constructed in this way. The approximation always provides a tight lower bound, is shown to be dramatically quicker to compute (with compute times reduced by up to four orders of magnitude), and gives remarkably accurate results for systems with large weights. Additionally, complementary datasets of weight systems satisfying the necessary but insufficient conditions for transversality were constructed, including considerations of the IP, reflexivity, and intradivisibility properties. Overall producing a classification of this weight system landscape, further confirmed with machine learning methods. Using the knowledge of this classification, and the properties of the presented approximation, a novel dataset of transverse weight systems consisting of 7 weights was generated for a sum of weights $\leq 200$; producing a new database of Calabi-Yau five-folds, with their respective topological properties computed. Further to this an equivalent database of candidate Calabi-Yau six-folds was generated with approximated Hodge numbers.
@@ 1 Introduction
Whilst Calabi-Yau three-folds are excellent candidates for superstring compactification from 10-dimensions, superstring theory also has interpretations within its parent theories of M-theory and F-theory, which are 11 and 12-dimensional, respectively. Therefore, to compactify these higher-dimensional theories down to 4-dimensions, higher-dimensional geometries are needed. M-theory compactification requires 7-dimensional $G2$-manifolds, with evermore elusive constructions, machine learning in this area has been initiated by recent work considering the related $G2$-structure geometries with success predicting their equivalent Hodge numbers Aggarwal:2023swe . Alternatively, F-theory compactification requires Calabi-Yau four-folds, where machine learning methods have been effective for the complete intersection construction [^1] We note recent successes in generating and machine learning Calabi-Yau five-folds Alawadhi:2023gxa , where the first systematic construction of complete intersection Calabi-Yau five-folds is presented, producing a large but inexhaustive list of new spaces. He:2020lbz ; Erbin:2021hmx , for which an exhaustive list has been determined Gray_2013 ; Gray_2014 ; however, machine learning methods have not yet been tested for the other constructions.
@@
This paper begins by detailing the Calabi-Yau construction of interest in §2, followed by analysis of the weight system data and respective topological invariants in §3, with detail on generated complementary datasets. In §4, the machine learning methods used are introduced, followed by their application, results, and interpretation. Central to this work, in §5, is the presentation of an approximation formula for computation of Hodge numbers of Calabi-Yau manifolds constructed via weighted projective spaces, providing a tight lower bound, and enormous improvements in compute time. In §6 this approximation, as well as the other related properties, is used to construct candidate transverse weight systems of 7 and 8 weights with sum of weights up to 200, along with the topological properties of the subsequent Calabi-Yau five- and six-folds respectively. Finally, in §7, results are summarised and outlook applications discussed.
@@ 2 Background
A polytope Bao:2021ofk is itself defined by a collection of $d$ hyperplane inequalities such that all points $x\in\mathbb{R}^{n}$ are in the polytope if $H\cdot x\geq b$ for some defining $d\times n$ matrix $H$ and constant $d$-vector $b$. The constituent parts of the polytope as defined by the intersection of the hyperplanes are the polytope faces, where 0-faces are vertices, 1-faces are edges, and so on up to $(n-1)$-faces which are facets. In the case where the vertices’ coordinates are all integers across the polytope, the polytope is a lattice polytope [^5] Lattice polytopes may also be physically interpreted as toric diagrams of quiver gauge theories Hanany:2005ve .. If the polytope contains a single lattice point in its strict interior, then the polytope is called interior point, or IP; due to the affine symmetry of the lattice this point can always be shifted to be the origin [^6] In the mathematics literature lattice polytopes with exclusively the origin in the strict interior are called Fano, since they lead to Fano varieties. Where the boundary lattice points are only the polytope’s vertices the polytope is terminal Fano, whilst where there’re extra boundary lattice points the polytope is canonical Fano Kasprzyk_2010 .. The respective toric fan for a lattice polytope is defined by constructing 1-cones which are lines connecting the origin to each vertex, then extending this line infinitely. The remaining higher cones of the fan are then defined by the intersections of the polytope hyperplanes [^7] To ensure a smooth toric variety the polytope can first be triangulated to resolve the singularities arising from the interior points of the facets.. The toric variety cox2011toric is then constructed from the fan through consideration of its respective dual fan, each cone in the fan has a dual cone which is the set of all points which inner product with points in the cone to produce a non-negative number. The union of the dual cones is the dual fan. Finally, the toric variety is defined as the maximal spectrum of the generators of the dual fan’s 1-cones, i.e. taking the dual fan 1-cone generators (vectors in $\mathbb{Z}^{n}$) and treating their entries as exponents of the coordinates in some $\mathbb{C}^{n}$, each generator provides a condition on the coordinate ring $\mathbb{C}^{n}$, and the resulting spectrum of maximal ideals of this quotient ring defines the toric variety.
@@
In this way, the weight systems of consideration for weighted projective spaces can be generalised to include combined weight systems (with many weight systems) for toric varieties. The hypersurface equation defining the potential Calabi-Yau in the weighted projective space then becomes a generic hypersurface in the toric varieties anticanonical divisor class Batyrev:1993oya ; with alternative interpretations as non-transverse hypersurfaces in weighted projective spaces Candelas:1994bu . Inverting the process of extracting weights from a polytope, polytopes can also be constructed from (combined) weight systems. However, before introducing this, the definition of a polytope’s dual is needed.
@@
Let us return to the construction of an IP polytope from an IP weight system. Here, the hyperplane equations defining the polytope include an equality for each weight system such that $\sum_{i}w_{i}x_{i}=w_{tot}$, and inequalities defined by $x_{i}\geq 0\ \forall i$ Batyrev:2020vnd . Through this construction, the point $x_{i}=1\ \forall i$ is manifestly contained within the polytope, since it naturally satisfies the equalities and inequalities; noting that an affine transformation can set this point to be the origin. This general polytope is hence always IP, however it may be rational and we are often more interested in its restriction to the integer lattice.
To be able to then define and check the IP property of a given weight system it suffices to construct the respective polytope, then restrict this polytope (with generally rational vertices) to the integer lattice. This restriction may slice parts of the rational polytope off (notably causing no change where the polytope is reflexive), and in doing this the new boundaries may intersect the origin. Therefore the restricted polytope may no longer contain the origin in its interior, and would thus not be IP Scholler:2018apc . In the cases where the origin does remains in the strict interior the respective lattice polytope is IP, and we define the weight system to be IP too.
All polytopes constructed from weight systems are simplices, since there are as many intersections of the single defining equality with the inequalities as there’re lattice dimensions (and also weights); equivalently, those from the larger combined weight systems are the union of simplices. For the weight system to be transverse there must be no further unavoidable singularities than the origin, and this translates to having no interior points on the polytope facets. It is where this occurs that weight systems can be IP but not transverse.
@@ 5 $IP(5)=R(5)\supset T(5)$ 184026 184026 7555
Table 2.1: Known relations between the sets of weight systems satisfying the IP ($IP(n)$), reflexive ($R(n)$), and transverse ($T(n)$) properties as the number of weights in the weight system increases. Additionally showing the sizes of the sets of weight systems satisfying those properties (denoted with ||$\cdot$|| ⋅ |) at each weight system size.
@@
then also allowing for the manifolds Euler number, $\chi$, to be computed from these Hodge numbers via $\chi=\sum_{k}(-1)^{k}b_{k}$.
@@ 3.1 The Four-folds Dataset
Figure 3.2: Scattered plots of the Hodge numbers of Calabi-Yau four-fold’s in weighted projective spaces. Plot (a) illustrates that the spaces are mirror symmetric to a high degree. Whilst (b) shows the relation between the two highest Hodge numbers, also compared with the constraint (2.10). Finally, the 3D plot illustrates the relation between the three independent Hodge numbers.
@@
This partitioning is discussed in more detail and put in a broader context in §3.2. For now, we just note that for highest weights larger than $\sim 5\times 10^{4}$, the $h^{1,1}$ values fall neatly into linear clusters. Motivated by the findings presented in Berman_2022 , we explore this behaviour of the dataset at hand with similar techniques. As it is evident from Figure 3.3, for large weights, there are eight peaks in the $h^{1,1}/w_{max}$ distribution (where the largest weight is the final weight in the weight system such that $w_{max}=w_{6}$). These are linear clusters in the $h^{1,1}$ vs $w_{max}$ plane, as shown in Figure 3(a). To neatly illustrate the clusters, we only considered systems with largest weights $w_{max}\gtrapprox 3\times 10^{5}$, which can be seen from Figure 3(b).
Figure 3.4: These plots focus on the clustering observed in Figure 3.3 for large weights, and show that the clustering analysis correctly reproduces the multi-linear behaviour.
@@ 3.2 Additional Weight Datasets
Table 3.1: The sizes of the subdatasets of 5-dimensional weight systems (i.e. 6 weights) in each part of the partition; along with the means and ranges across all weight values in each subdataset.
@@ 3.3 Principal Component Analysis
Linear behaviour in distributions can analysed through principal component analysis (PCA). This unsupervised machine learning technique extracts an orthonormal basis for the dataset in question, with basis vectors ranked according to their degree of contribution towards the variance in the data’s distribution. The basis is computed as eigenvectors of the dataset’s covariance matrix, where the symmetric nature of the matrix ensures real eigenvalues that can be ordered decreasingly and then used to rank the basis. The normalised eigenvalues are named the explained variance, and provide a measure of relative importance of each eigenvector (the larger explained variance the more important the respective eigenvector). For a prespecified desired degree of representation, a dataset can be projected onto the first $i$ eigenvectors in the ranked basis such that the sum of the respective first $i$ normalised eigenvalues exceeds the desired proportion of representation. In this sense, PCA is often used as a dimensionality reduction technique.
@@
given to the nearest integer. Particularly displaying similar lower bounds throughout, whilst higher mean and maximum values for non-reflexive subdatasets, and substantially larger ranges for the CY data.
@@
The similarity between the R and nR subdatasets PCA projections indicates classification architectures will likely find identification of this property harder. Whilst the higher skewed values of PCA projection, as well as the far larger maximum values, for the CY datasets will perhaps be used by the architectures to aid learning.
@@ 4 Machine Learning
For the prediction of cohomological data, which has a very wide range of possible values, a NN regressor was used. Since the input data are small (just six integers), a simple architecture with few layers was enough for this problem. Specifically, we used the built-in multi-layer perception regressor from scikit-learn, with the following features: $(16,32,16)$ layer structure, ReLU activation, $MSE$ loss, and Adam optimiser. We chose a training-test split of $80-20$, and performed a 5-fold cross-validation for each investigation. The batch size was set to $200$, and we imposed an upper bound of $250$ epochs (the network could stop before that if it reached convergence). Regarding the performance measures, we focused on the following three:
@@ 4.1 Regressing Hodge Numbers
We again see that the accuracy drops from left to right, according to the chosen subset. Finally, for completeness, we also present our results on $h^{1,2}$ and $\chi$. Since both of them can be zero, the MAPE measure does not apply to these cases, and therefore we omit it.
@@
The fact that higher cohomologies in Calabi-Yau four-folds are harder to learn with neural networks has already appeared in the literature, in He:2020lbz . Although they analysed a different construction of four-folds, i.e. complete intersection Calabi-Yau’s (CICY), their results also indicate that $h^{1,1}$ is the only Hodge number that can be successfully learnt to high levels of precision with a fully connected networks. Convolutional neural networks have proven to be the most efficient ones for learning Hodge numbers in the case of CICY’s, which are described by a matrix; however, this kind of architecture is not suitable for Calabi-Yau’s in weighted projective spaces, where the input consists of a one-dimensional vector.
@@ 4.1.1 NN Gradient Saliency
In the saliency analysis performed here, each NN is differentiated with respect to each of the inputs, and the differential evaluated at each of the test data inputs. The absolute values of these gradient components are then averaged over the test dataset, as well as averaged over the run repetitions – here repeating the investigation with randomised 80:20 train:test splits for 100 independent NNs of the same architecture. Since function scales can vary through the NN layers, the relative saliency values are the features of interest; they are represented, for the neural network predicting $h^{1,1}$, in Figure 4.1. The 6 weights of the input weight systems are represented by 6 boxes, where lighter colours indicate higher saliency values and larger relative importance. These results show that the NNs focus on the weights in each system according to their size. They prioritise the information encoded in the lower weights, while the largest weights seem not to play an as important role. This implies that the networks are not exploiting the clustering behaviour shown in Figure 3.4, previously discussed. To be expected if we consider that the vast majority of weights actually lies in the “bulk” of such scattered plot, while the linear behaviour is only evident for systems with extremely large weights.

### v2

@@ Abstract
Calabi-Yau four-folds may be constructed as hypersurfaces in weighted projective spaces of complex dimension 5 defined via weight systems of 6 weights. In this work, neural networks were implemented to learn the Calabi-Yau Hodge numbers from the weight systems, where gradient saliency and symbolic regression then inspired a truncation of the Landau-Ginzburg model formula for the Hodge numbers of any dimensional Calabi-Yau constructed in this way. The approximation always provides a tight lower bound, is shown to be dramatically quicker to compute (with computation times reduced by up to four orders of magnitude), and gives remarkably accurate results for systems with large weights. Additionally, complementary datasets of weight systems satisfying the necessary but insufficient conditions for transversality were constructed, including considerations of the interior point, reflexivity, and intradivisibility properties. Overall producing a classification of this weight system landscape, further confirmed with machine learning methods. Using the knowledge of this classification, and the properties of the presented approximation, a novel dataset of transverse weight systems consisting of 7 weights was generated for a sum of weights $\leq 200$; producing a new database of Calabi-Yau five-folds, with their respective topological properties computed. Further to this an equivalent database of candidate Calabi-Yau six-folds was generated with approximated Hodge numbers.
@@ 1 Introduction
Whilst Calabi-Yau three-folds are excellent candidates for superstring compactification from 10-dimensions, superstring theory also has interpretations within its parent theories of M-theory and F-theory, which are 11 and 12-dimensional, respectively. Therefore, to compactify these higher-dimensional theories down to 4-dimensions, higher-dimensional geometries are needed. M-theory compactification requires 7-dimensional manifolds Gherardini:2023uyx , notably G2222-manifolds exhibiting evermore elusive constructions, machine learning in this area has been initiated by recent work considering the related G2222-structure geometries with success predicting their equivalent Hodge numbers Aggarwal:2023swe . Alternatively, F-theory compactification requires Calabi-Yau four-folds, where machine learning methods have been effective for the complete intersection construction [^1] We note recent successes in generating and machine learning Calabi-Yau five-folds https://doi.org/10.1002/prop.202300262 , where the first systematic construction of complete intersection Calabi-Yau five-folds is presented, producing a large but inexhaustive list of new spaces. He:2020lbz ; Erbin:2021hmx , for which an exhaustive list has been determined Gray_2013 ; Gray_2014 ; however, machine learning methods have not yet been tested for the other constructions.
@@
This paper begins by detailing the Calabi-Yau construction of interest in §2, followed by analysis of the weight system data and respective topological invariants in §3, with detail on generated complementary datasets. In §4, the machine learning methods used are introduced, followed by their application, results, and interpretation. Central to this work, in §5, is the presentation of an approximation formula for computation of Hodge numbers of Calabi-Yau manifolds constructed via weighted projective spaces, providing a tight lower bound, and enormous improvements in computation time. In §6 this approximation, as well as the other related properties, is used to construct candidate transverse weight systems of 7 and 8 weights with sum of weights up to 200, along with the topological properties of the subsequent Calabi-Yau five- and six-folds respectively. Finally, in §7, results are summarised and outlook applications discussed.
@@ 2 Background
A polytope Bao:2021ofk is itself defined by a collection of $d$ hyperplane inequalities such that all points $x\in\mathbb{R}^{n}$ are in the polytope if $H\cdot x\geq b$ for some defining $d\times n$ matrix $H$ and constant $d$-vector $b$. The constituent parts of the polytope as defined by the intersection of the hyperplanes are the polytope faces, where 0-faces are vertices, 1-faces are edges, and so on up to $(n-1)$-faces which are facets. In the case where the vertices’ coordinates are all integers across the polytope, the polytope is a lattice polytope [^5] Lattice polytopes can also be physically interpreted as toric diagrams of quiver gauge theories Hanany:2005ve .. If the polytope contains a single lattice point in its strict interior, then the polytope is called interior point, or IP; due to the affine symmetry of the lattice this point can always be shifted to be the origin [^6] In the mathematics literature lattice polytopes with exclusively the origin in the strict interior are called Fano, since they lead to Fano varieties. Where the boundary lattice points are only the polytope’s vertices the polytope is terminal Fano, whilst where there’re extra boundary lattice points the polytope is canonical Fano Kasprzyk_2010 .. The respective toric fan for a lattice polytope is defined by constructing 1-cones which are lines connecting the origin to each vertex, then extending each line infinitely. The remaining higher cones of the fan are then defined by the intersections of the polytope hyperplanes [^7] To ensure a smooth toric variety the polytope can first be triangulated to resolve the singularities arising from the interior points of the facets.. The toric variety cox2011toric is then constructed from the fan through consideration of its respective dual fan, each cone in the fan has a dual cone which is the set of all points whose inner product with points in the cone produces a non-negative number. The union of all dual cones is the dual fan. Finally, the toric variety is defined as the maximal spectrum of the generators of the dual fan’s 1-cones, i.e. taking the dual fan 1-cone generators (vectors in $\mathbb{Z}^{n}$) and treating their entries as exponents of the coordinates in some $\mathbb{C}^{n}$, each generator provides a condition on the coordinate ring $\mathbb{C}^{n}$, and the resulting spectrum of maximal ideals of this quotient ring defines the toric variety.
@@
In this way, the weight systems of consideration for weighted projective spaces can be generalised to include combined weight systems (with many weight systems) for toric varieties. The hypersurface equation defining the potential Calabi-Yau in the weighted projective space then becomes a generic hypersurface in the toric variety’s anticanonical divisor class Batyrev:1993oya ; with alternative interpretations as non-transverse hypersurfaces in weighted projective spaces Candelas:1994bu . Inverting the process of extracting weights from a polytope, polytopes can also be constructed from (combined) weight systems. However, before introducing this, the definition of a polytope’s dual is needed.
@@
Let us return to the construction of an IP polytope from an IP weight system. Here, the hyperplane equations defining the polytope include an equality for each weight system such that $\sum_{i}w_{i}x_{i}=w_{tot}$, and inequalities defined by $x_{i}\geq 0\ \forall i$ Batyrev:2020vnd . Through this construction, the point $x_{i}=1\ \forall i$ is manifestly contained within the polytope, since it naturally satisfies the equalities and inequalities; noting that an affine transformation can set this point to be the origin. This general polytope is hence always IP, however it may be rational and we are often more interested in its restriction to a lattice.
To be able to then define and check the IP property of a given weight system it suffices to construct the respective polytope, and consider it as existing on the crudest lattice it can (that generated by the polytopes vertices). The dual polytope can then be generated from this, and respectively the dual lattice (all real points which dot product with all points in the polytope’s lattice to integers), however the vertices of the dual polytope may not lie on the dual lattice, and thus a restriction is required by taking the convex hull of dual lattice points that lie within/on this dual polytope. This restriction may slice parts of the dual polytope off, producing a smaller dual polytope, which when taking the dual again will produce an new version of the original polytope, which we define to be the integer polytope of interest, and in doing this the new boundaries may intersect the origin. Therefore the new restricted polytope may no longer contain the origin in its interior, and would thus not be IP Scholler:2018apc . In the cases where the origin does remains in the strict interior the respective lattice polytope is IP, and we define the weight system to be IP too.
All real polytopes constructed from weight systems are simplices, since there are as many intersections of the single defining equality with the inequalities as there’re lattice dimensions (and also weights); equivalently, those from the larger combined weight systems are the union of simplices. However the restriction to the relevant lattice as described above may generalise the polytope causing the lattice polytopes to be unions of simplices also. For the weight system to be transverse there must be no further unavoidable singularities than the origin, and this translates to having no interior points on the polytope facets. It is where this occurs that weight systems can be IP but not transverse.
@@ 5 $IP(5)=R(5)\supset T(5)$ 184026 184026 7555
Table 2.1: Known relations between the sets of weight systems satisfying the IP ($IP(n)$), reflexive ($R(n)$), and transverse ($T(n)$) properties as the number of weights in the weight system increases. Additionally showing the sizes of the sets of weight systems satisfying those properties (denoted with ||$\cdot$|| ⋅ |) at each weight system size. The weight systems for dimensions $\geq 7$ have not been fully computed yet.
@@
then also allowing for the manifolds’ Euler number, $\chi$, to be computed from these Hodge numbers via $\chi=\sum_{k}(-1)^{k}b_{k}$.
@@ 3.1 The Four-folds Dataset
Figure 3.2: Scattered plots of the Hodge numbers of Calabi-Yau four-fold’s in weighted projective spaces. Plot (a) illustrates that the spaces are mirror symmetric to a high degree. Whilst (b) shows the relation between the two highest Hodge numbers, also compared with the constraint (2.10). Finally, the 3D plot (c) illustrates the relation between the three independent Hodge numbers.
@@
This partitioning is discussed in more detail and put in a broader context in §3.2. For now, we just note that for highest weights larger than $\sim 5\times 10^{4}$, the $h^{1,1}$ values fall neatly into linear clusters. Motivated by the findings presented in Berman_2022 , we explore this behaviour of the dataset at hand with similar techniques. As it is evident from Figure 3.3, for large weights, there are eight peaks in the $h^{1,1}/w_{max}$ distribution (where the largest weight is the final weight in the weight system, such that $w_{max}=w_{5}$). These are linear clusters in the $h^{1,1}$ vs $w_{max}$ plane, as shown in Figure 4(a). To neatly illustrate the clusters, we only considered systems with largest weights $w_{max}\gtrapprox 3\times 10^{5}$, which can be seen from Figure 4(b).
Figure 3.4: These plots focus on the clustering observed in Figure 3.3 for large weights, and show that the clustering analysis correctly reproduces the multi-linear behaviour. The clusters are shown as peaks in the $h_{1,1}/w_{max}$ histogram in (a) and as lines in the $h_{1,1}$ vs $w_{max}$ plane in (b).
@@ 3.2 Additional Weight Datasets
Table 3.1: The sizes of the subdatasets of weight systems of 6 weights in each part of the partition; along with the means and ranges across all weight values in each subdataset.
@@ 3.3 Principal Component Analysis
Linear behaviour in distributions can be analysed through principal component analysis (PCA). This unsupervised machine learning technique extracts an orthonormal basis for the dataset in question, with basis vectors ranked according to their degree of contribution towards the variance in the data’s distribution. The basis is computed as eigenvectors of the dataset’s covariance matrix, where the symmetric nature of the matrix ensures real eigenvalues that can be ordered decreasingly and then used to rank the basis. The normalised eigenvalues are named the explained variance, and provide a measure of relative importance of each eigenvector (the larger the explained variance the more important the respective eigenvector). For a prespecified desired degree of representation, a dataset can be projected onto the first $i$ eigenvectors in the ranked basis such that the sum of the respective first $i$ normalised eigenvalues exceeds the desired proportion of representation. In this sense, PCA is often used as a dimensionality reduction technique.
@@
given to the nearest integer. They display similar lower bounds throughout, whilst higher mean and maximum values for non-reflexive subdatasets, and substantially larger ranges for the CY data.
@@
The similarity between the R and nR subdatasets’ PCA projections indicates classification architectures will likely find identification of this property harder. Whilst the higher skewed values of PCA projection, as well as the far larger maximum values, for the CY datasets will perhaps be used by the architectures to aid learning.
@@ 4 Machine Learning
For the prediction of cohomological data, which has a very wide range of possible values, a NN regressor was used. Since the input data are small (just six integers), a simple architecture with few layers was enough for this problem. Specifically, we used the built-in multi-layer perception regressor from scikit-learn, with the following features: $(16,32,16)$ layer structure, ReLU activation, $MSE$ loss, and Adam optimiser. We chose a training-test split of $80:20$, and performed a 5-fold cross-validation for each investigation. The batch size was set to $200$, and we imposed an upper bound of $250$ epochs (the network could stop before that if it reached convergence). Regarding the performance measures, we focused on the following three:
@@ 4.1 Regressing Hodge Numbers
We again see that the accuracy drops from left to right, according to the chosen subset. Finally, for completeness, we also present our results on $h^{1,2}$ and $\chi$ in Table 4.4. Since both of them can be zero, the MAPE measure does not apply to these cases, and therefore we omit it.
@@
The fact that higher cohomologies in Calabi-Yau four-folds are harder to learn with neural networks has already appeared in the literature, in He:2020lbz . Although they analysed a different construction of four-folds, i.e. complete intersection Calabi-Yau’s (CICY), their results also indicate that $h^{1,1}$ is the only Hodge number that can be successfully learnt to high levels of precision with fully connected networks. Convolutional neural network variants have exhibited the highest accuracies on the CICY matrix inputs Erbin:2020srm , however due to the permutation symmetry of the configuration matrices, as well as the weight system vectors for the construction considered here, the benefits of the convolutional architecture’s focus on local properties is lost, we therefore stick to the more general dense feed-forward architectures.
@@ 4.1.1 NN Gradient Saliency
In the saliency analysis performed here, each NN is differentiated with respect to each of the inputs, and the differential evaluated at each of the test data inputs. The absolute values of these gradient components are then averaged over the test dataset, as well as averaged over the run repetitions – here repeating the investigation with randomised 80:20 train:test splits for 100 independent NNs of the same architecture. Since function scales can vary through the NN layers, the relative saliency values are the features of interest; they are represented, for the NNs predicting $h^{1,1}$, in Figure 4.1. The 6 weights of the input weight systems are represented by 6 boxes, where lighter colours indicate higher saliency values and larger relative importance. These results show that the NNs focus on the weights in each system according to their size. They prioritise the information encoded in the lower weights, while the largest weights seem not to play an as important role. This implies that the networks are not exploiting the clustering behaviour shown in Figure 3.4, previously discussed. Perhaps, to be expected if we consider that the vast majority of weights actually lie in the “bulk” of the scatter plots, while the linear behaviour is only evident for systems with extremely large weights.